# Phase error rate estimation in QKD with imperfect detectors

### v2

@@ IV.2.2 Key Generation Rounds after basis-independent Filter
When basis-efficiency mismatch is present, one must figure out the phase error rate in the key generation rounds, which are filtered using the $Z$ basis. However the rounds for testing are filtered using the $X$ basis. These filtering steps are not identical. Therefore it becomes very difficult to prove rigorous bounds on the phase error rate based on the observed data. One of the main contributions of this work is a rigorous derivation of such bounds, without relying on asymptotic behaviour or IID assumptions.
@@ IV.3 Cost of removing the basis-independent loss assumption
In removing the basis-independent loss assumption from phase error estimation, we will need to define metrics $\delta_{1},\delta_{2}$, which will quantify the deviation from ideal behaviour. We will now explain how these metrics are defined.
@@ IV.4 Sampling
    $e^{\text{obs}}_{X}$ is the fraction of the $n_{X}$ testing rounds that resulted in the $G^{\text{con,F}}_{(XX),(\neq)}$ outcome. We have access to $e^{\text{obs}}_{X}$ in the protocol, since it is something we actually observe. (Recall that this is the only error rate we observe).
@@
    $\Pr(\bm{\tilde{e}^{\text{key}}_{XX}}\geq\bm{\tilde{e}^{\text{obs}}_{XX}}+\gamma_{\text{serf}})_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}}$ $\leq e^{-\gamma_{\text{serf}}^{2}f_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})}.$ (21)
    We define
    $\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})\coloneq\sqrt{\frac{\ln(1/\varepsilon_{\text{AT-a}}^{2})}{f_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})}}\implies e^{-\left(\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})\right)^{2}f_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})}=\varepsilon_{\text{AT-a}}^{2}.$ (22)
@@
  • $\bm{\tilde{e}^{\text{key}}_{XX}}\leftrightarrow\bm{\tilde{e}^{\text{key}}_{ZX}}$: We utilize the definition of $\delta_{1}$ stated in Section IV.3. Since the POVM elements generating $\bm{\tilde{e}^{\text{key}}_{ZX}}$ ($G^{\text{con,F}}_{(ZX),(\neq)}$) and $\bm{\tilde{e}^{\text{key}}_{XX}}$ ($G^{\text{con,F}}_{(XX),(\neq)}$) are close, we expect the bounds obtained on $\bm{\tilde{e}^{\text{key}}_{ZX}}$ and $\bm{\tilde{e}^{\text{key}}_{XX}}$ to also be close. This is made precise in the following lemma proved in Section C.2.
@@
    Recall that $\bm{n_{K}}$ is obtained by discarding rounds from $\bm{\tilde{n}_{K}}$ based on $\{{F}_{(Z),(\text{con})},\mathrm{I}-{F}_{(Z),(\text{con})}\}$ measurements. We will essentially show that very few rounds are discard in this step, using Eq. 19. To do so, we prove the following Lemma in Appendix C.
@@
Thus we have relationships Eqs. 20, 23, 28 and 31 between all the error rates that hold with high probability. These can all be combined using straightforward but cumbersome algebra (see Appendix D), to obtain
@@
Using the above expression requires us to know the values of $\tilde{n}_{K}$ and $\tilde{n}_{X}$ which we do not. This problem is easily resolved by noting all the $\gamma$s are decreasing functions of $\tilde{n}_{K}$ and $\tilde{n}_{X}$, and that $\tilde{n}_{K}(\tilde{n}_{X})$ cannot be larger than $n_{K}(n_{X})$ (since we discard rounds to from the former to the latter) . Thus, we can replace $\tilde{n}_{K}$ with $\bm{n_{K}}$ and $\tilde{n}_{X}$ with $\bm{n_{X}}$ and obtain
@@
Let us investigate the behaviour of Eq. 34 in the limit $\delta_{1},\delta_{2}\rightarrow 0$. Recall that $\gamma_{\text{bin}}^{\varepsilon_{\text{AT}}}(n,\delta)$ was defined as the value of $c$ such that $F(n,\delta,c)=\sum_{i=n(\delta+c)}^{n}{n\choose k}\delta^{i}(1-\delta)^{n-i}\leq\varepsilon_{\text{AT}}^{2}$. However, notice that $\delta\rightarrow 0\implies F(n,\delta,c)\rightarrow 0$ for any value of $c$. Therefore, $\delta\rightarrow 0\implies\gamma_{\text{bin}}(n,\delta)\rightarrow 0$. Setting these limits in Eq. 34, we recover the result Eq. 14 for the case where the basis-independent loss assumption is satisfied.
@@
$\delta_{1}$ $=2\norm{\sqrt{F^{\text{key}}_{\text{con}}}G^{\text{key}}_{\neq}\sqrt{F^{\text{key}}_{\text{con}}}-\sqrt{F^{\text{test}}}G^{\text{test}}_{\neq}\sqrt{F^{\text{test}}}}_{\infty},$
@@ IV.5 Relation to other work on basis-efficiency mismatch
We will now discuss several prior works on addressing basis-efficiency mismatch in the literature. In a broad sense, the reformulation of the detector setup as first implementing a filtering step followed by further measurements is an important ingredient in many of these works (although the precise details may differ). However, all of them perform an IID asymptotic analysis, where there is no need for finite-size sampling arguments we use (such as Lemmas 2 and 3). Instead one can directly associate the various error rates with POVM measurements on a single round state, and the analysis is greatly simplified.
@@
Ref. [12] proposed the first security proof of QKD in the presence of Eve controlled basis-efficiency mismatch, in the asymptotic regime for IID collective attacks. It required the assumption that a qubit is received on Bob’s side, for which it required the use of detector-decoy methods. Furthermore it argued that the basis-efficiency mismatch can be removed entirely (for qubits received by Bob) if one randomly swaps the 00 and the $1$ detector. Note that this trick only works for the qubit subspace and does not hold for higher photon numbers (see Section VI.3). Ref. [11] extended this work without having to assume qubit detections, but still performs an IID asymptotic analysis. In Ref. [38], the numerical framework for key rate computations was proposed and used to compute IID asymptotic key rates for perfectly characterized and fixed basis-efficiency mismatch. In Ref. [13] the numerical framework [38] was used to compute IID asymptotic key rates for a toy model where Eve was allowed to induce basis efficiency mismatch via the use of two spatio-temporal modes. Recently, Ref. [14] improved upon [12] by obtaining tighter estimates on the phase error rate in the presence of basis efficiency mismatch, but again assumed IID collective attacks in the asymptotic regime, and single qubits received on Bob’s side. The assumption of single qubits received by Bob was later removed in the follow up work [15]. In both [14, 15], Eve is not allowed to control the efficiency mismatch. Finally, MDI-QKD [10] addresses all detector side-channels in the finite-size regime against coherent attacks, but requires a drastically different protocol implementation compared to standard QKD.
@@ V Application to Decoy-state BB84
First, we rigorously prove the security for the variable-length decoy-state BB84 protocol in the entropic uncertainty relations framework (note that Ref. [30] actually implicitly implements a variable-length protocol). In fact [30] considers a protocol with iterative sifting  [^5] This has been formulated in a variety of ways in the literature. In general, we use this phrase for protocols that have a “loop” phase, where some actions are taken repeatedly until certain conditions are met. where the total number of rounds in the protocol is not fixed a priori, and depends on observations made during the protocol, which are announced in a round-by-round manner. However a justification for this step is not provided. This is important because certain kinds of iterative sifting can lead to subtle issues in the security analysis (see Ref. [45] for some issues and Ref. [46] for solutions). We do not fix this problem directly in this work. Instead, we consider a protocol with a fixed total number of signals sent, which avoids this problem. Second, we make rigorous certain technical steps in [30] (regarding entropic calculations on states conditioned on events), which we point out where applicable (see Remark 19). Third, our phase error estimates do not require the assumption of basis-independent loss unlike that of [30]. We also avoid a particular Taylor series approximation used by [30] ([47, Eq. 22]), and therefore our phase error estimation procedure yields a true bound without any approximations. Finally we also clarify certain aspects of the decoy analysis undertaken in [30]. In particular, we careful differentiate between random variables and an observed value of the random variable, and also properly condition on relevant events in our presentation. We stress that while we clarify and make rigorous certain steps in [30], our main contribution in this work is the variable-length security proof of decoy-state BB84 in the presence of detector imperfections.
@@
Recently, a more accessible version of the security proof in Ref. [30] was written in Ref. [48]. We note that while Ref. [48] is certainly quite pedagogical, it does not address all the issues described above.
@@ V.3 Decoy Analysis
The application of Hoeffdings inequality here is subtle, and is made rigorous in Lemmas 12 and 13 in Appendix E (see also Ref. [43]). Note that in general, the photon numbers of every pulse in the protocol are chosen independently, since Alice chooses intensity independently for each pulse. However, here we are interested in photon numbers corresponding to rounds that led to a specific outcome $O$. Since we postselect pulses based on the outcome, we can no longer claim that the photon numbers of these pulses (pulses that led to outcome $O$) are sampled independently, or that intensities of these pulses are chosen independently. This is because they now depend on Eve’s attack. Rather, Lemmas 12 and 13 rely on exploiting the fact that conditioned on any fixed sequence of photon numbers of the pulses, the intensities are chosen independently of one another. One can therefore apply Hoeffdings inequality. Then, since the resulting statements holds for any fixed sequence of photon numbers, the conditioning on this event can be removed.
@@
Let $\mathcal{S}_{\text{constraints}}$ denote the set of equations inside the probability in the above expressions. Therefore we have $\Pr(\mathcal{S}_{\text{constraints}})\geq 1-9\varepsilon_{\text{AT-d}}^{2}$.
@@ V.4 Bounds on zero and one photon statistics
For any event $O\in\{X,X_{\neq},K\}$, the relevant bounds on the zero-photon and single-photon components can be obtained by algebraic manipulation of the expressions in $\mathcal{S}_{\text{constraints}}$. In general, any method for bounding the relevant zero-photon and single-photon components using $\mathcal{S}_{\text{constraints}}$ suffices. In this work, we follow exactly the steps taken by Ref. [30, Appendix A] to obtain these bounds. Thus, we only write the final expressions here. We define
@@
The lower bound on the zero-photon component is given by [30, Eq. 2]
@@
The lower bound on the one-photon component is given by [30, Eq. 3]
@@
$\left(\bm{n_{O,\mu_{2}}}^{-}-\bm{n_{O,\mu_{3}}}^{+}-\frac{\mu_{2}^{2}-\mu_{3}^{2}}{\mu_{1}^{2}}\left(\bm{n_{O,{1}}}^{+}-\mathcal{B}^{\text{decoy}}_{\text{min}-0}(\bm{n_{O,\mu_{\vec{k}}}})/\tau_{0}\right)\right).$
The upper bound on the one-photon component is given by [30, Eq. 4]
@@
Since $\Pr(\mathcal{S}_{\text{constraints}})\geq 1-9\varepsilon_{\text{AT-d}}^{2}$, and Eqs. 49, 50 and 51 follow from the expressions in $\mathcal{S}_{\text{constraints}}$. We have
@@ VI Results
We will now apply our results to a decoy-state BB84 protocol with realistic detectors. To do so, we start by outlining a recipe for using this work to compute key rates in Section VI.1. We will then specify the canonical model for our detectors with efficiency mismatch in Section VI.2. We will apply the recipe to our model in Section VI.3. Finally, we will plot the key rate we obtain in Section VI.4.

### v3

@@ IV.2.2 Key Generation Rounds after basis-independent Filter
When basis-efficiency mismatch is present, one must figure out the phase error rate in the key generation rounds, which are filtered using the $Z$ basis. However the rounds for testing are filtered using the $X$ basis. These filtering steps are not identical. Therefore it becomes very difficult to prove rigorous bounds on the phase error rate based on the observed data. One of the main contributions of this work is a rigorous derivation of such bounds, without relying on asymptotic behavior or IID assumptions.
@@ IV.3 Cost of removing the basis-independent loss assumption
In removing the basis-independent loss assumption from phase error estimation, we will need to define metrics $\delta_{1},\delta_{2}$, which will quantify the deviation from ideal behavior. We will now explain how these metrics are defined.
@@ IV.4 Sampling
    $e^{\text{obs}}_{X}$ is the fraction of the $n_{X}$ testing rounds that resulted in the $G^{\text{con,F}}_{(XX),(\neq)}$ outcome. We have access to $e^{\text{obs}}_{X}$ in the protocol, since it is something we actually observe.
@@
    $\Pr(\bm{\tilde{e}^{\text{key}}_{XX}}\geq\bm{\tilde{e}^{\text{obs}}_{XX}}+\gamma_{\text{serf}})_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}}$ $\leq e^{-2\gamma_{\text{serf}}^{2}f_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})}.$ (21)
    Using the definition from Eq. 13, we have
    $\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})=\sqrt{\frac{\ln(1/\varepsilon_{\text{AT-a}}^{2})}{2f_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})}}\implies e^{-\left(\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})\right)^{2}2f_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})}=\varepsilon_{\text{AT-a}}^{2}.$ (22)
@@
  • $\bm{\tilde{e}^{\text{key}}_{XX}}\leftrightarrow\bm{\tilde{e}^{\text{key}}_{ZX}}$: We utilize the definition of $\delta_{1}$ stated in Section VI.2. Since the POVM elements generating $\bm{\tilde{e}^{\text{key}}_{ZX}}$ ($G^{\text{con,F}}_{(ZX),(\neq)}$) and $\bm{\tilde{e}^{\text{key}}_{XX}}$ ($G^{\text{con,F}}_{(XX),(\neq)}$) are close, we expect the bounds obtained on $\bm{\tilde{e}^{\text{key}}_{ZX}}$ and $\bm{\tilde{e}^{\text{key}}_{XX}}$ to also be close. This is made precise in the following lemma proved in Section C.2.
@@
    Recall that $\bm{n_{K}}$ is obtained by discarding rounds from $\bm{\tilde{n}_{K}}$ based on $\{{F}_{(Z),(\text{con})},\mathrm{I}-{F}_{(Z),(\text{con})}\}$ measurements. We will essentially show that very few rounds are discarded in this step, using Eq. 19. To do so, we prove the following Lemma in Appendix C.
@@
Thus we have relationships Eqs. 20, 23, 28 and 31 between all the error rates, whose complements hold with high probability. These can all be combined using straightforward but cumbersome algebra (see Appendix D), to obtain
@@
Using the above expression requires us to know the values of $\tilde{n}_{K}$ and $\tilde{n}_{X}$ which we do not. This problem is easily resolved by noting all the $\gamma$s are decreasing functions of $\tilde{n}_{K}$ and $\tilde{n}_{X}$, and that $\tilde{n}_{K}(\tilde{n}_{X})$ cannot be smaller than $n_{K}(n_{X})$ (since we discard rounds to from the former to the latter) . Thus, we can replace $\tilde{n}_{K}$ with $\bm{n_{K}}$ and $\tilde{n}_{X}$ with $\bm{n_{X}}$ and obtain
@@
Let us investigate the behavior of Eq. 34 in the limit $\delta_{1},\delta_{2}\rightarrow 0$. Recall that $\gamma_{\text{bin}}^{\varepsilon_{\text{AT}}}(n,\delta)$ was defined as the value of $c$ such that $F(n,\delta,c)=\sum_{i=n(\delta+c)}^{n}{n\choose k}\delta^{i}(1-\delta)^{n-i}\leq\varepsilon_{\text{AT}}^{2}$. However, notice that $\delta\rightarrow 0\implies F(n,\delta,c)\rightarrow 0$ for any value of $c$. Therefore, $\delta\rightarrow 0\implies\gamma_{\text{bin}}(n,\delta)\rightarrow 0$. Setting these limits in Eq. 34, we recover the result Eq. 14 for the case where the basis-independent loss assumption is satisfied.
@@
$\delta_{1}$ $=2\norm{\sqrt{F^{\text{key}}_{\text{con}}}G^{\text{key}}_{\neq}\sqrt{F^{\text{key}}_{\text{con}}}-\sqrt{F^{\text{test}}_{\text{con}}}G^{\text{test}}_{\neq}\sqrt{F^{\text{test}}_{\text{con}}}}_{\infty},$
@@ IV.5 Relation to other work on basis-efficiency mismatch
Paper Coherent Attacks Finite-size Eve has (bounded) control over detectors Hardware Modifications Notes
Fung et al. [12] Yes No Yes Detector Decoy [43] (To remove Qubit assumption) -
Lydersen et al. [11] Yes No Yes None Handles wide range of multi-mode models
Øystein Marøy et al. [33] Yes No Yes None Handles wide range of multi-mode models
Winick et al. [44] No No No None -
Zhang et al. [13] No No Yes* None * Only two modes
Bochkov et al. [14] No No No None Qubit assumption on Bob
Trushechkin et al. [15] No No No None -
Grasselli et al. [45] No No No Detector decoy [43], requires tunable beam splitter Does not require detector characterization
Marcomini et al. [34] Yes No No No Qubit assumption Bob, can handle some source imperfections.
This work Yes Yes Yes None
Table 2: Comparison of prior work on phase error rate estimation in the presence of basis efficiency mismatch. Note that Ref. [13] relies on numerical evidence for a part of the proof (bounding weight outside the subspace of low photon numbers).
@@
We will now discuss several prior works on addressing basis-efficiency mismatch in the literature (see Table 2). In a broad sense, the reformulation of the detector setup as first implementing a filtering step followed by further measurements is an important ingredient in many of these works (although the precise details may differ). However, all of them perform an asymptotic analysis, where there is no need for finite-size sampling arguments we use (such as Lemmas 2 and 3). Instead one can directly associate the various error rates with POVM measurements on a single round state, and the analysis is greatly simplified.
@@
Ref. [12] proposed the first security proof of QKD in the presence of (bounded) Eve controlled basis-efficiency mismatch, in the asymptotic regime. It required the assumption that a qubit is received on Bob’s side, for which it required the use of detector-decoy methods. Furthermore it also argued that the basis-efficiency mismatch can be removed entirely (for qubits received by Bob) if one randomly swaps the 00 and the $1$ detector. Note that this trick only works for the qubit subspace and does not hold for higher photon numbers (see Section VI.4). Refs. [11, 33] extended this work without having to assume qubit detections, but still performs an asymptotic analysis. In Ref. [44], the numerical framework for key rate computations was proposed and used to compute IID asymptotic key rates for perfectly characterized and fixed basis-efficiency mismatch. In Ref. [13] the numerical framework [44] was used to compute IID asymptotic key rates for a toy model where Eve was allowed to induce basis efficiency mismatch via the use of two spatio-temporal modes. Recently, Ref. [14] improved upon [12] by obtaining tighter estimates on the phase error rate in the presence of basis efficiency mismatch, but again assumed IID collective attacks in the asymptotic regime, and single qubits received on Bob’s side. The assumption of single qubits received by Bob was later removed in the follow up work [15]. In both [14, 15], Eve is not allowed to control the efficiency mismatch. A recent work [45] again considers a scenario with IID collective attacks in the asymptotic regime, but has the advantage of not requiring prior characterization of the detector parameters. Another recent work [34] combines qubit flaws in the source with efficiency mismatch in the detectors for coherent attacks, but is valid only in the asymptotic regime, and that Bob always receives a qubit. Finally, MDI-QKD [10] addresses all detector side-channels in the finite-size regime against coherent attacks, but requires a drastically different protocol implementation compared to standard QKD.
@@ V Application to Decoy-state BB84
First, we rigorously prove the security for the variable-length decoy-state BB84 protocol in the entropic uncertainty relations framework (note that Ref. [35] actually implicitly implements a variable-length protocol). In fact [35] considers a protocol with iterative sifting  [^6] This has been formulated in a variety of ways in the literature. In general, we use this phrase for protocols that have a loop phase, where some actions are taken repeatedly until certain conditions are met. where the total number of rounds in the protocol is not fixed a priori, and depends on observations made during the protocol, which are announced in a round-by-round manner. However a justification for this step is not provided. This is important because certain kinds of iterative sifting can lead to subtle issues in the security analysis (see Ref. [52] for some issues and Ref. [53] for solutions). We do not fix this problem directly in this work. Instead, we consider a protocol with a fixed total number of signals sent, which avoids this problem. Second, we make rigorous certain technical steps in [35] (regarding entropic calculations on states conditioned on events), which we point out where applicable (see Remark 21). Third, our phase error estimates do not require the assumption of basis-independent loss unlike that of [35]. We also avoid a particular Taylor series approximation used by [35] ([54, Eq. 22]), and therefore our phase error estimation procedure yields a true bound without any approximations. Finally we also clarify certain aspects of the decoy analysis undertaken in [35]. In particular, we careful differentiate between random variables and an observed value of the random variable, and also properly condition on relevant events in our presentation. We stress that while we clarify and make rigorous certain steps in [35], our main contribution in this work is the variable-length security proof of decoy-state BB84 in the presence of detector imperfections.
@@
Recently, a more accessible version of the security proof in Ref. [35] was written in Ref. [55]. While Ref. [55][Version 2] addresses many of the above concerns for fixed-length protocols, it does not deal with detector imperfections or variable-length protocols.
@@ V.3 Decoy Analysis
The application of Hoeffdings inequality here is subtle, and is made rigorous in Lemmas 13 and 14 in Appendix E (see also Ref. [50]). Note that in general, the photon numbers of every pulse in the protocol are chosen independently, since Alice chooses intensity independently for each pulse. However, here we are interested in photon numbers corresponding to rounds that led to a specific outcome $O$. Since we postselect pulses based on the outcome, we can no longer claim that the photon numbers of these pulses (pulses that led to outcome $O$) are sampled independently, or that intensities of these pulses are chosen independently. This is because they now depend on Eve’s attack. Rather, Lemmas 13 and 14 rely on exploiting the fact that conditioned on any fixed sequence of photon numbers of the pulses, the intensities are chosen independently of one another. One can therefore apply Hoeffdings inequality. Then, since the resulting statements holds for any fixed sequence of photon numbers, the conditioning on this event can be removed.
@@
Let $\mathcal{S}_{\text{constraints}}$ denote the set of inequalities inside the probability in the above expressions. Therefore we have $\Pr(\mathcal{S}_{\text{constraints}})\geq 1-9\varepsilon_{\text{AT-d}}^{2}$.
@@ V.4 Bounds on zero and one photon statistics
For any event $O\in\{X,X_{\neq},K\}$, the relevant bounds on the zero-photon and single-photon components can be obtained by algebraic manipulation of the expressions in $\mathcal{S}_{\text{constraints}}$. In general, any method for bounding the relevant zero-photon and single-photon components using $\mathcal{S}_{\text{constraints}}$ suffices. In this work, we follow exactly the steps taken by Ref. [35, Appendix A] to obtain these bounds. Thus, we only write the final expressions here. We define
@@
The lower bound on the zero-photon component is given by [35, Eq. 2]
@@
The lower bound on the one-photon component is given by [35, Eq. 3]
@@
$\left(\bm{n_{O,\mu_{2}}}^{-}-\bm{n_{O,\mu_{3}}}^{+}-\frac{\mu_{2}^{2}-\mu_{3}^{2}}{\mu_{1}^{2}}\left(\bm{n_{O,\mu_{1}}}^{+}-\mathcal{B}^{\text{decoy}}_{\text{min}-0}(\bm{n_{O,\mu_{\vec{k}}}})/\tau_{0}\right)\right).$
The upper bound on the one-photon component is given by [35, Eq. 4]
@@
Since $\Pr(\mathcal{S}_{\text{constraints}})\geq 1-9\varepsilon_{\text{AT-d}}^{2}$, and Eqs. 49, 50 and 51 follow from the expressions in $\mathcal{S}_{\text{constraints}}$, we obtain
@@ VI Results
We will now apply our results to a decoy-state BB84 protocol with realistic detectors. To do so, we start by outlining a recipe for using this work to compute key rates in Section VI.1. We will then specify the canonical model for our detectors with efficiency mismatch in Section VI.3. We will apply the recipe to our model in Section VI.4. Finally, we will plot the key rate we obtain in Section VI.5.